OpenStudy (anonymous):

(1 +cos 2 theta)/(2 cos theta) = cos theta

OpenStudy (anonymous):

Verify that is an identity...Can someone help?

OpenStudy (anonymous):

\[\frac{ 1+\cos 2\theta }{2cos \theta } \] =\[\frac{ 2\cos ^{2}\theta }{ 2\cos \theta } =\cos \theta \]

OpenStudy (anonymous):

Thank You

OpenStudy (anonymous):

\[\tan 157.5\] Use half angle?

OpenStudy (anonymous):

\[\tan 157.5=\tan \left( 180-22.5 \right)=-\tan 22.5\] \[\tan 2x=\frac{ 2\tan x }{ 1-\tan ^{2}x },substitute 2x=45,\tan 2x=\tan 45=1\] x=22.5 \[1=\frac{ 2\tan x }{ 1-\tan ^{2} x}\] \[1-\tan ^{2}x=2\tan x\] \[\tan ^{2}x+2\tan x -1=0 \] \[\tan x=\frac{ -2\pm \sqrt{\left( -2 \right)^{2}-4\times1\times-1} }{2\times1 }\] \[\tan x=\frac{ -2\pm \sqrt{4+4} }{ 2 }\] \[\tan x=\frac{ -2\pm \sqrt{4\times2} }{ 2 }\] \[\tan x=\frac{ -2\pm2\sqrt{2} }{ 2 }\] \[\tan x=-1\pm \sqrt{2}\] x=22.5 lies in first quadrant, he\[\tan 22.5=-1+\sqrt{2}\]nce tan x is positive\[\tan 157.5=-\tan 22.5=1-\sqrt{2}\]